# Drinfeld singular moduli, hyperbolas, units

### v1

@@ Abstract.
The authors were supported by the IRN MaDeF (CNRS), IRN GandA (CNRS), and by ANR-20-CE40-0003 Jinvariant. They also benefitted from support of Laboratoire de Mathématiques Nicolas Oresme (Caen) and Laboratoire de Mathématiques de Besançon. The final steps of this work where performed while the last author was in residence at Institut Mittag-Leffler in Djursholm, Sweden, during the semester in Analytic Number Theory: tack så mycket! We dedicate this work to Bas Edixhoven and to Andrea Surroca, who passed away way too early.
March 31, 2024
@@ 4.4. Upper bound for $h(\alpha)$
$\mathop{\mathrm{Card}}C_{\varepsilon}(D)$ $\leq\max_{\begin{subarray}{c}a\in A_{+}\\ |a|=|D|^{1/2}\end{subarray}}\!\!2^{\omega(a)}\,\,\Bigl{(}\!\!\!\sum_{\begin{subarray}{c}a\in A_{+}\\ |a|=|D|^{1/2}\\ |\frac{\sqrt{D}}{2e}-a|<\varepsilon|a|\end{subarray}}\max\{1,q\varepsilon|\gcd\nolimits_{2}(a,D)|\}\Bigr{)}$
@@ 4.5. Lower bound for $h(\alpha)$
Let $K$ be an imaginary quadratic extension of $k$, and let $\alpha\in\mathbb{C}_{\infty}$ be a singular modulus of discriminant $D$. Let $D_{K}$ be the disciminant of $\mathcal{O}_{K}$. If $\alpha$ has CM by an order of conductor $f$, we have $D=D_{K}f^{2}$ and as soon as $|D_{K}|\geq q^{4}$ we obtain
$h(\alpha)\geq$ $\quad\frac{\log_{q}|D_{K}|}{10}\left(\frac{1}{2}-\frac{1}{\sqrt{q}+1}\right)\ln q-\left[\frac{5q-3}{4q-4}+\frac{1}{2}-\frac{1}{\sqrt{q}+1}+\frac{8}{\ln q}\right]\frac{\ln q}{5}$
@@
Note that $\alpha\not=0$ since $|D|>1$. Let $\Phi$ be the Drinfeld module of rank 2 defined by
@@
We now invoke the results of Fu-Tsun Wei, valid when $\Phi$ has CM by $\mathcal{O}_{K}$. Apply Theorem 5.3 page 887 of [Wei20], Lemma 5.5 page 887 of [Wei20], and equation (5.5) page 889 of [Wei20] to obtain
@@
where $g_{K}$ is the genus of $K$ and $q_{K}$ is the cardinality of the constant field of $K$, similarly $g_{k}$ is the genus of $k$, and $\gamma_{K}$ is the Euler-Kronecker constant of $K$, and $q_{\infty}=\#\mathbb{F}_{\infty}$. We know $g_{k}=0$, $q_{\infty}\geq q$, $q_{K}\geq q$, but we need an estimate on $\gamma_{K}$. We will use the work of Ihara [Iha06], who gives in (1.4.6) page 420 the inequality
@@
(28) $h_{\mathrm{Tag}}(\Phi)\geq\frac{g_{K}-1}{2}\ln q-\ln q+\frac{1}{2}\ln q-\frac{g_{K}}{\sqrt{q}+1}\ln{q}-\frac{q+1}{4(q-1)}\ln{q},$
hence
(29) $h_{\mathrm{Tag}}(\Phi)\geq g_{K}\left(\frac{1}{2}-\frac{1}{\sqrt{q}+1}\right)\ln q-\frac{5q-3}{4q-4}\ln{q}$
@@
(30) $h(\Phi)\geq\frac{g_{K}}{5}\left(\frac{1}{2}-\frac{1}{\sqrt{q}+1}\right)\ln q-\frac{5q-3}{20q-20}\ln{q}-\frac{6}{5}.$
By Riemann-Hurwitz (Theorem 7.16 page 90 of [Ros02]) we get $g_{K}\geq\frac{1}{2}\log_{q}|D_{K}|-1$, which leads to:
(31) $h(\Phi)\geq\frac{\log_{q}|D_{K}|}{10}\left(\frac{1}{2}-\frac{1}{\sqrt{q}+1}\right)\ln q-\left[\frac{5q-3}{4q-4}+\frac{1}{2}-\frac{1}{\sqrt{q}+1}+\frac{6}{\ln q}\right]\frac{\ln q}{5},$
which implies the claim in the case where $\Phi$ has CM by the maximal order $\mathcal{O}_{K}$.
@@
Combining (23), (25), (32), and direct estimates on (34) lead to the claim for general orders. Let us compute a valid explicit constant by bounding the right hand side of (34) from above. We start with the particular case $\deg(f)=1$, where we get
(35) $\frac{1}{2}\sum_{v|f}\mathrm{deg}(v)e_{f}(v)\leq\frac{q}{(q-1)^{2}}.$
@@
(38) $\frac{1}{2}\sum_{v|f}\mathrm{deg}(v)e_{f}(v)\leq\frac{4q^{2}}{(q-1)^{2}}\log_{q}^{+}\log_{q}|f|.$
@@
(39) $h_{\mathrm{Tag}}(\Phi)\geq h_{\mathrm{Tag}}(\Phi^{\prime})+\frac{1}{2}\log_{q}|f|-\frac{4q^{2}}{(q-1)^{2}}\log_{q}^{+}\log_{q}|f|-\frac{q}{(q-1)^{2}},$
and finally in (31)
$h(\Phi)\geq$ $\frac{\log_{q}|D_{K}|}{10}\left(\frac{1}{2}-\frac{1}{\sqrt{q}+1}\right)\ln q-\left[\frac{5q-3}{4q-4}+\frac{1}{2}-\frac{1}{\sqrt{q}+1}+\frac{6}{\ln q}\right]\frac{\ln q}{5}$
@@
which gives the claim. ∎
@@
Let $\alpha$ be a singular modulus of discriminant $D=D_{K}f^{2}$. If $|D_{K}|\geq q^{4}$, then
@@
(41) $-\left[\frac{5q-3}{4q-4}+\frac{1}{2}-\frac{1}{\sqrt{q}+1}+\frac{8}{\ln q}\right]\frac{\ln q}{5}\geq-3\ln q.$
@@
The following proposition is due to Ran and gives a formula which is the Drinfeld module analogue of the Nakkajima-Taguchi [NT91] formula for CM elliptic curves. We recall it here for completeness.
@@
Let $\Phi$ be a Drinfeld module of rank 2 that has CM by $\mathcal{O}=\mathbb{F}_{q}[T]+f\mathcal{O}_{K}$, order of conductor $f\in\mathbb{F}_{q}[T]$. Then $\Phi$ is isogenous to another rank 2 Drinfeld module $\Phi^{\prime}$ with CM by the maximal order $\mathcal{O}_{K}$, and we have

### v2

@@ Abstract.
The authors were supported by the IRN MaDeF (CNRS), IRN GandA (CNRS), and by ANR-20-CE40-0003 Jinvariant. They also benefitted from support of Laboratoire de Mathématiques Nicolas Oresme (Caen) and Laboratoire de Mathématiques de Besançon. The final steps of this work where performed while the last author was in residence at Institut Mittag-Leffler in Djursholm, Sweden, during the semester in Analytic Number Theory: tack så mycket! We thank Tim Trudgian for interesting feedback. We dedicate this work to Bas Edixhoven and to Andrea Surroca, who passed away way too early.
April 11, 2024
@@ 4.4. Upper bound for $h(\alpha)$
(20) $\mathop{\mathrm{Card}}C_{\varepsilon}(D)$ $\leq\max_{\begin{subarray}{c}a\in A_{+}\\ |a|=|D|^{1/2}\end{subarray}}\!\!2^{\omega(a)}\,\,\Bigl{(}\!\!\!\sum_{\begin{subarray}{c}a\in A_{+}\\ |a|=|D|^{1/2}\\ |\frac{\sqrt{D}}{2e}-a|<\varepsilon|a|\end{subarray}}\max\{1,q\varepsilon|\gcd\nolimits_{2}(a,D)|\}\Bigr{)}$
@@ 4.5. Lower bound for $h(\alpha)$
In this section, we obtain two different lower bounds for the height of a singular modulus, given by Proposition 4.17 and Corollary 4.19.
The first lower bound is easy:
@@
The second lower bound will follow from the following proposition:
@@
Let $K$ be a separable, imaginary quadratic extension of $k$, and let $\alpha\in\mathbb{C}_{\infty}$ be a singular modulus having CM by an order of $\mathcal{O}_{K}$ of conductor $f$. Then we have:
$h(\alpha)\geq$ $\quad\frac{\log_{q}|D_{K}|}{10}\left(\frac{1}{2}-\frac{1}{\sqrt{q}+1}\right)\ln q-\left[\frac{7q-5}{4q-4}+\frac{8}{\ln q}\right]\frac{\ln q}{5}$
@@
If $\alpha=0$ then $K=\mathbb{F}_{q^{2}}(T)$, $|D_{K}|=1$ and $f=1$. Since $h(0)=0$, the lower bound of the proposition holds in that case.
From now on we assume that $\alpha\not=0$. Let $\Phi$ be the Drinfeld module of rank 2 defined by
@@
Consider first the special case where $\Phi$ has CM by $\mathcal{O}_{K}$. We then invoke the results of Fu-Tsun Wei. Apply Theorem 5.3 page 887 of [Wei20], Lemma 5.5 page 887 of [Wei20], and equation (5.5) page 889 of [Wei20] to obtain
@@
where $g_{K}$ is the genus of $K$, $q_{K}$ is the cardinality of the constant field of $K$, $g_{k}$ is the genus of $k$, $\gamma_{K}$ is the Euler-Kronecker constant of $K$ and $q_{\infty}=\#\mathbb{F}_{\infty}$. We need an estimate on $\gamma_{K}$. We will use the work of Ihara [Iha06], who gives in (1.4.6) page 420 the inequality
@@
(37) $h_{\mathrm{Tag}}(\Phi)\geq\frac{g_{K}-1}{2}\ln q_{K}-\ln q+\frac{1}{2}\ln q-\frac{g_{K}}{\sqrt{q}+1}\ln{q}-\frac{q+1}{4(q-1)}\ln{q}.$
We distinguish now two cases.
Suppose first that $K\not=\mathbb{F}_{q^{2}}(T)$. Then $q_{K}=q$ and (37) yields
$h_{\mathrm{Tag}}(\Phi)\geq g_{K}\left(\frac{1}{2}-\frac{1}{\sqrt{q}+1}\right)\ln q-\frac{5q-3}{4q-4}\ln{q}.$
By the Riemann-Hurwitz theorem (Theorem 7.16 of [Ros02]) we have $g_{K}\geq\frac{1}{2}\log_{q}|D_{K}|-1$, hence we get:
(38) $h_{\mathrm{Tag}}(\Phi)\geq\frac{\log_{q}|D_{K}|}{2}\left(\frac{1}{2}-\frac{1}{\sqrt{q}+1}\right)\ln q-\left(\frac{7q-5}{4q-4}-\frac{1}{\sqrt{q}+1}\right)\ln q.$
Suppose now that $K=\mathbb{F}_{q^{2}}(T)$. Then $q_{K}=q^{2}$ and $g_{K}=0$, hence (37) gives
(39) $h_{\mathrm{Tag}}(\Phi)\geq-\frac{7q-5}{4q-4}\ln q.$
Since $|D_{K}|=1$ in that case, we see from (38) and (39) that we have, in both cases:
(40) $h_{\mathrm{Tag}}(\Phi)\geq\frac{\log_{q}|D_{K}|}{2}\left(\frac{1}{2}-\frac{1}{\sqrt{q}+1}\right)\ln q-\frac{7q-5}{4q-4}\ln q.$
@@
Combining (33), (40), (41), and direct estimates on (42) leads to the claim for general orders. Let us compute a valid explicit constant by bounding the right hand side of (42) from above. We start with the particular case $\deg(f)=1$, where we get
$\frac{1}{2}\sum_{v|f}\mathrm{deg}(v)e_{f}(v)\leq\frac{q}{(q-1)^{2}}.$
@@
$\frac{1}{2}\sum_{v|f}\mathrm{deg}(v)e_{f}(v)\leq\frac{4q^{2}}{(q-1)^{2}}\log_{q}^{+}\log_{q}|f|.$
@@
$h_{\mathrm{Tag}}(\Phi)\geq h_{\mathrm{Tag}}(\Phi^{\prime})+\frac{1}{2}\log_{q}|f|-\frac{4q^{2}}{(q-1)^{2}}\log_{q}^{+}\log_{q}|f|-\frac{q}{(q-1)^{2}},$
and finally, using (33) and (40):
$h(\alpha)\geq$ $\frac{\log_{q}|D_{K}|}{10}\left(\frac{1}{2}-\frac{1}{\sqrt{q}+1}\right)\ln q-\left[\frac{7q-5}{4q-4}+\frac{6}{\ln q}\right]\frac{\ln q}{5}$
@@
which gives the proposition. ∎
We can now state our second lower bound for the height of a singular modulus.
@@
Let $K$ be a separable, imaginary quadratic extension of $k$, and let $\alpha\in\mathbb{C}_{\infty}$ be a singular modulus having CM by an order of $\mathcal{O}_{K}$ of discriminant $D$. Then:
@@
(46) $-\left[\frac{7q-5}{4q-4}+\frac{8}{\ln q}\right]\frac{\ln q}{5}\geq-3\ln q.$
@@
In the proof of Proposition 4.18 we have used the following proposition. It is due to Ran and gives a formula that is the Drinfeld module analogue of the Nakkajima-Taguchi [NT91] formula for CM elliptic curves. We recall it here for completeness.
@@
Let $\Phi$ be a Drinfeld module of rank 2 that has CM by $\mathcal{O}=\mathbb{F}_{q}[T]+f\mathcal{O}_{K}$, order of conductor $f\in\mathbb{F}_{q}[T]$. Then $\Phi$ is isogenous to another rank 2 Drinfeld module Φsuperscriptnormal-$\Phi^{\prime}$ with CM by the maximal order $\mathcal{O}_{K}$, and we have
@@
4.6. Proof of Theorem 1.2
In this section, we use the bounds obtained in Sections 4.4 and 4.5 to finally prove Theorem 1.2.